# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

### v1

@@ 3.3 Two-step Verdict Classification (TVC)
Three-Class Classification: The first step involves categorizing each claim into one of three categories: Supported, Refuted, or Not Enough Information (NEI). Given a claim-evidence pair $(C,E)$, we employ a classification function:
$P_{\text{3}}=f_{\text{Three-Class}}(C,E),$ (1)
where
$P_{\text{3}}=\left[P_{\text{SUP}},P_{\text{REF}},P_{\text{NEI}}\right]$
represents the predicted probabilities. Here, $f_{\text{Three-Class}}$ denotes a neural network model based on an advanced BERT-based architecture, trained to classify each claim into one of the three categories. The model leverages a fine-tuned BERT-derived approach and is optimized using Cross Entropy Loss, ensuring that the predicted probability distribution closely aligns with the true labels.
Binary Classification: If the three-class model does not predict NEI, an additional classification step is applied to further refine the decision between Supported and Refuted. In this step, we use another classification function:
$P_{\text{2}}=f_{\text{Binary-Class}}(C,E),$ (2)
where
$P_{\text{2}}=\left[P_{\text{SUP}},P_{\text{REF}}\right].$
Here, $f_{\text{Binary-Class}}$ is a neural network based on an advanced BERT-based architecture, specifically trained to distinguish between Supported and Refuted claims. Unlike the three-class model, this binary classifier leverages a fine-tuned BERT-derived model and is trained using Focal Loss  Lin et al. (2018) to mitigate class imbalance and improve performance on challenging examples.
Finally, the predicted class $\hat{y}$ is determined based on the following decision rule:
$\hat{y}=\begin{cases}\text{NEI},&\text{if }\hat{y}_{\text{3}}=\text{NEI}\\ \hat{y}_{\text{2}},&\text{if }P_{\text{2}}>P_{\text{3}}\text{ and }\hat{y}_{\text{3}}\neq\hat{y}_{\text{2}}\\ \hat{y}_{\text{3}},&\text{if }P_{\text{3}}\geq P_{\text{2}}\text{ or }\hat{y}_{\text{3}}=\hat{y}_{\text{2}}.\end{cases}$ (3)
This method effectively balances the strengths of both classification models, ensuring improved accuracy for complex claims while maintaining high performance on straightforward cases.
@@ 4.2 Experimental Setup
Additionally, the large language model was fine-tuned on a distributed setup with 4×H100 GPUs, utilizing a structured prompt reformulation process. The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning. This ensured effective task-specific learning while aligning with pre-trained knowledge. Further details on training setup, prompt design, and preprocessing are in Appendix B.

### v2

@@ 3.3 Two-step Verdict Classification (TVC)
3.4 Three-Class Classification
The initial step involves classifying each claim-evidence pair $(C,E)$ into one of three categories: Supported, Refuted, or Not Enough Information (NEI). We define a three-class classification function as follows:
$\mathbf{P}_{3}=f_{\text{3-class}}(C,E),$ (1)
where $\mathbf{P}_{3}=[P^{\text{NEI}}_{3},P^{\text{SUP}}_{3},P^{\text{REF}}_{3}]$ denotes the predicted probability distribution over the three classes, and $f_{\text{3-class}}$ represents a neural network model based on an advanced BERT-derived architecture, fine-tuned for this task. The predicted label is determined by:
$\hat{y}_{3}=\arg\max_{k\in\{\text{NEI},\text{SUP},\text{REF}\}}P^{k}_{3},$ (2)
where $\hat{y}_{3}\in\{\text{NEI},\text{SUP},\text{REF}\}$ is the class with the highest predicted probability. The model is optimized using Cross-Entropy Loss, ensuring that $\mathbf{P}_{3}$ aligns closely with the ground-truth distribution.
3.5 Binary Classification
When the three-class model does not predict NEI (i.e., $\hat{y}_{3}\neq\text{NEI}$), a secondary binary classification step refines the decision between Supported and Refuted. We define a binary classification function:
$\mathbf{P}_{2}=f_{\text{2-class}}(C,E),$ (3)
where $\mathbf{P}_{2}=[P^{\text{SUP}}_{2},P^{\text{REF}}_{2}]$ represents the predicted probability distribution over the two classes, and $f_{\text{2-class}}$ is a neural network based on a fine-tuned BERT-derived architecture, specifically trained for binary classification. The predicted label is given by:
$\hat{y}_{2}=\arg\max_{k\in\{\text{SUP},\text{REF}\}}P^{k}_{2},$ (4)
where $\hat{y}_{2}\in\{\text{SUP},\text{REF}\}$ corresponds to the class with the highest probability. This model is optimized using Focal Loss Lin et al. (2018) to address class imbalance and enhance performance on difficult examples.
3.6 Final Decision Rule
The final predicted class, denoted $\hat{y}$, is derived by integrating the outputs of the three-class and binary classification models. When the three-class model predicts $\hat{y}_{3}=\text{NEI}$, the decision is immediate. In all other cases, a comparison is made between the maximum probability from the three-class model (excluding the NEI class) and that from the binary model. This process is formalized by the following decision rule:
$\hat{y}=\begin{cases}\text{NEI},&\text{if }\hat{y}_{3}=\text{NEI},\\ \hat{y}_{3},&\text{if }P^{\text{max}}_{3}>P^{\text{max}}_{2},\\ \hat{y}_{2},&\text{otherwise},\end{cases}$ (5)
where $P^{\text{max}}_{3}=\max(P^{\text{SUP}}_{3},P^{\text{REF}}_{3})$ represents the highest probability assigned by the three-class model to either the Supported or Refuted class, and $P^{\text{max}}_{2}=\max(P^{\text{SUP}}_{2},P^{\text{REF}}_{2})$ denotes the corresponding maximum probability from the binary model.
This approach leverages the complementary strengths of both models: the three-class model provides a broad categorization, including the critical NEI class, while the binary model offers refined discrimination between Supported and Refuted claims. By prioritizing the three-class prediction when its confidence exceeds that of the binary model, the method ensures robust handling of complex claims while maintaining high accuracy for well-defined cases.
@@ 4.2 Experimental Setup
Additionally, the large language model was fine-tuned on a distributed setup with A100 GPUs, utilizing a structured prompt reformulation process. The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning. This ensured effective task-specific learning while aligning with pre-trained knowledge. Further details on training setup, prompt design, and preprocessing are in Appendix B.